OpenStudy (rebecca1588):

Which equation shows the substitution method being used to solve the system of linear equations? x – y = 7 x = y + 3 A. (y + 3) – y = 7 B. <<<<<<<

OpenStudy (anonymous):

I think it is C, I am unsure.

OpenStudy (rebecca1588):

Thanks :)

OpenStudy (anonymous):

This is confusing... x-3=y and x+7=y

OpenStudy (anonymous):

oh wait nvm x-3=y... x-7=y...

OpenStudy (rebecca1588):

If I get this 1 wrong its no big deal ^^

OpenStudy (anonymous):

D cannot be right because x-x would equal zero not seven

OpenStudy (anonymous):

A could be right, it shows the first equation... x-y=7

OpenStudy (anonymous):

I believe it is A